# Formation and Study of a Spherical Plasma Liner for Plasma-Jet-Driven Magneto-Inertial Fusion

### 20 Jan 2024 v1

@@ Abstract
Plasma-jet-driven magneto-inertial fusion (PJMIF) is an alternative approach to controlled nuclear fusion which aims to realize a line-replaceable dense plasma liner as a repetitive spherical compression driver. First experimental measurements on the formation of a spherical Argon plasma liner formed from 36 discrete pulsed plasma jets are obtained on the Plasma Liner Experiment (PLX). Liner uniformity and morphology, plasma density, and temperature are assessed at times throughout the implosion process.
I Introduction:
@@
Magneto-inertial fusion (MIF, also called magnetized target fusion or MTF) is an approach to controlled nuclear fusion which marries aspects of both magnetic confinement fusion and inertial confinement fusion methods, enabling controlled fusion at more modest values of both confinement time and density [1, 2]. Many MIF compression approaches and configurations have been proposed and studied, including the MagLIF Z-pinch configuration [3], shear-flow-stabilized Z pinches [4], spherical piston-driven liquid metal liners [5], rotationally stabilized liquid metal liners[6], and staged-gas-puff Z-pinches[7]. Notably, MagLIF experiments have successfully produced thermonuclear conditions and significant neutron yields from the compression of preheated magnetized deuterium fuel, on implosion timescales significantly longer than can be tolerated in pure inertial fusion.
@@
The plasma-jet-driven magneto-inertial fusion (PJMIF) [8] concept is an MIF approach in which the requisite heavy compressive liner is formed from the merging and spherical convergence of an array of discrete plasma jets. PJMIF aims to access an attractive subsection of the MIF physics parameters space, namely that of the moderate velocity (50-150 km/s) compression of high-beta high-Hall-parameter targets which have shown promise in MagLIF [2]. If successful, this architecture would enable repetitive line-replaceable operation with a significant standoff distance between the fusion plasma and driver hardware, which is a significant challenge facing all pulsed fusion concepts, in particular those seeking to utilize electrical pulsed-power as the primary energy source [1, 8].
@@
There are several key questions regarding the viability of the PJMIF approach, including whether conditions of sufficient energy density and fusion gain can be attained while maintaining the physical standoff that is the attractiveness of the approach, whether sufficient uniformity of a spherical liner can be attained from the merger of discrete jets in the face of jet-on-jet shock waves and deceleration-phase Rayleigh-Taylor instabilities [9, 10, 11], and whether effective magnetic insulation of a standoff target plasma can be achieved in spherical implosion geometry. Prior investigations have shed light on several of these questions, with analytical and computational modeling predicting the feasibility of significant fusion gain [12], the realization of high-performance modular plasma guns to produce controlled plasma jets [13], and laboratory experiments studying the merging and formation processes of multiple plasma jets at representative velocities [14, 15].
@@
In the present work, fully spherical plasma liners have been formed and investigated at the Plasma Liner Experiment (PLX). This experiment uses 36 individual argon plasma jets, enabling observation of the integrated liner assembly process for the first time. The discussion of the full liner experimental campaign is organized as follows: Section II details the pulsed power setup and plasma jet formation and tuning, Section III describes what plasma properties are measured and the diagnostics used to do so, and Section IV presents the results of the diagnostic suite and implications for PJMIF as a whole. Considerations necessary for continuation and an eventual reactor-scale device are explored in Section VI as well as a concluding summary of points.
II Experiment:
@@
PLX consists of a 3-meter diameter chamber and 36 attached plasma guns, shown in Fig. 1 along with a rendered cross section of the chamber and notional plasma liner and target within [16]. The plasma guns are positioned quasi-spherically over the chamber, with a typical full-angle between adjacent guns of 22 degrees. Chamber vacuum base pressure just before the plasma is injected into the chamber is typically around 1 torr, increasing to 10 Torr when a turbopump gate valve is closed before the plasma shot.
@@
The plasma guns that supply the liner-forming plasma jets are contoured-gap coaxial plasma railguns - the unique contoured shape of the inner electrode alleviates the blowby instability in such geometries that can hinder outgoing jet quality. The guns used in the present PLX work are similar to those described in detail in Ref[13], although the PLX ones are a newer and more compact design. The stored energy per gun per pulse is 4.5-6 kJ, which accelerates jets of total mass 1 mg of argon plasma to typical tunable velocities between 30 and 70 km/s out of a barrel that is 8.5 cm in diameter. Most of the plasma gun capacitor banks and spark-gap switches are mounted directly onto the chamber to minimize transfer inductance to the plasma load and increase energy coupling. Due to space access limitations around the lowest ports of the PLX chamber, nine guns are instead linked to their capacitor banks via an array of transmission lines, referred to hereafter as transmission line (TL) guns. Plasma properties of the jets at the time of ejection have been explored in previous publications [17]; typically for a given jet around ejection time, eV and cm-3.
@@
Four distinct collections of capacitor banks are activated when firing a PLX shot. The first is a high voltage (19.0 kV) modified-Blumlein trigger network for the plasma gun gas valves (GV). This sends 36 trigger pulses down coaxial cables trigger 36 spark gap switches operating the gas valve capacitor banks of the plasma guns. Each jet has a separate 100 J GV capacitor bank, powering a custom fast-acting gas puff valve that supplies the neutral argon gas injection into the guns. An adjustable time delay passes while gas valves operate and neutral gas flows into the bore of the plasma guns, typically 480 microseconds in the current experiments. Thirdly, another modified-Blumlein trigger network fires, to breakdown the parallel spark-gaps for the main gun banks. These banks have 575 F rated to 5 kV peak voltage, and provide most of the current to form and accelerate the plasma jet. The main bank is discharged across a set of six spark-gap switches to the coaxial rail electrodes, again to minimize transfer inductance and prolong switch electrode lifetime. Operating voltages are set to 4.0 kV for non-TL guns, while the TL guns described above are charged 4.5 kV to compensate for the increased transfer inductance of their transmission lines. Time is defined as when the main trigger pulse for the main accelerating banks is sent. Pressure adjustments can be made to many of the spark-gap switches on an individual basis to ensure they all break down properly with the trigger pulse. A resistive circuit in parallel with the PFN spark-gap trigger switches produces pre-ionization at the breech - this preionization reduces time jitter of the bulk plasma breakdown and subsequent acceleration generated by the main bank.
@@
These component voltages can be adjusted for varying effects: a higher main trigger voltage can reduce firing jitter but increases probability of prefiring and arcing between components; main bank charge voltage establishes the overall energy with which the jets are ionized and accelerated; GV charge voltages establish how much gas is injected into the breach and subsequently accelerated. Tuning the individual jet velocity is essential in this experiment, in order to ensure the jets arrive in the chamber at the same time and merge together with as high degree of spherical symmetry as possible. Adjusting the GV bank voltage provided the most robust means of tuning jet velocity, with the effect that a constant main bank energy accelerating a lower gas injected to higher ultimate velocity. To achieve a nominal jet front velocity of 50 km/s for each of the 36 jets, adjustments to this GV bank voltage for each individual gun are made approximately 2 seconds to firing to fine-tune the respective jet velocity. This tuning method enables significantly more flexible and easier adjustment than the previous method of introducing ballast impedence wires in series with the GV described in Appendix B of Ref[18]. Results in this paper are obtained using a nominal best tune of each gun to 50 km/s. The process of determining the best tune is described in Section III.2. Mach number , with , where adiabatic index is assumed and is the Argon ion mass. To estimate M, we will use calculations from previous work examining the exiting jet for temperature and . With these values, estimated ejection Mach number in the present experiment is around .
@@
GV and main bank discharges are monitored using a set of Rogowski coils (six for the main bank, one for GV). Jet ejection time is taken as 2.5 s after the peak of the main bank signal. This ejection time relative to the signal peak is due to the Rogowski coils being positioned at the transmission lines at the back of the gun before the plasma travels down the barrel length and is expelled which takes about an additional 2.5 s. Fig. 2 shows a set of all 36 main bank Rogowski signals, the mean signal and standard deviation zone around it, and the mean ejection time given by the dashed line. Current ringing that can be observed after the primary accelerating current corresponding to the signal peak, as the underdamped capacitor circuit discharges its remaining energy into residual trailing plasma in the gun.
III Diagnostics:
@@
A suite of diagnostics is used to interrogate the plasma liner properties throughout the merge and stagnation process. Time-resolved laser interferometry measurements of line-integrated plasma density are acquired over the duration of each shot, and a multi-frame fast camera images the plasma at several times throughout each shot. A collection of single-frame wide-view cameras are also used for imaging. Photodiode time-of-flight pairs are used measure individual jet velocities for each of the 36 plasma jets. Individual gun performance is also monitored using Rogowski coils that measure the main accelerating current of each gun (see Fig. 2). An imaging spectrometer provides electron temperature and charge state information and high-resolution spectrometer is used to verify electron density and jet slowing at chamber center. Diagnostics lacking inherent time-resolution are instead triggered at varying times for different shots to build up an average picture of temporal evolution of the liner.
@@ III.1 Imaging
III.2 Photodiodes
Each gun is equipped with two light-collecting fibers positioned along the jet trajectory and spaced by 2 cm which feed to two photodiodes having sufficiently fast risetime to track the jet (5 ns). The 2 cm separation distance was selected as a compromise between two competing considerations: there must be enough time separation between the signals that the available digitization rate of 40 MHz does not significantly contribute to uncertainty (i.e. there are sufficient samples separating the two signals), and the closer together the fibers are the signals are more likely to be closely correlated in shape which enables easier analysis.
Fig. 4 shows a diagram of the two photodiode views of a jet, at two times of jet emergence from the gun cavity. The corresponding locations on normalized sample signal readouts are also indicated, along with the region within which a cross correlation is applied. The bounds of the cross correlation region are selected to be a few microseconds before the signal rises, and a few microseconds after the signal peak. This provides the jet front velocity. The jets consist of a distribution of directed velocities with the trailing portion being slower than the jet front, resulting in an extended column of plasma. The velocity analysis focused on the front region is used because it is this front forms the spherical liner to compress the MIF target and is therefore most relevant. The photodiode analysis process is automated and repeated for each gun and enables velocities to be measured immediately following each shot and tuned as desired by adjusting the GV bank voltage - this we call the tuned velocity setting.
This photodiode pair time-of-flight velocimetry methodology was cross-checked by taking a set of shots with a single gun at a time while positioning a single-tip Langmuir probe at chamber center, and obtaining a jet velocity via the plasma time-of-flight to the probe () and comparing to the velocity inferred via the photodiode signals (). A typical comparison is shown in Fig. 5. We define the jet arrival time to the probe as the location corresponding to 20% of the full integral of the probe signal. This generally matches well with the first peak-like feature in the signal, as seen in subfigure b.
@@ III.2 Photodiodes
III.3 High-Resolution Monochromator
@@ III.4 Multi-Chord Interferometry
Line-integrated electron density is obtained in this work via a multi-chord heterodyne laser interferometry system [20]. The laser emission is produced using a 320 mW, 561 nm solid state laser, is split into multiple lower-power beams and injected into the chamber through five single-mode fiber optic cables with each probe beam about 0.3 cm in diameter. The scene chords of the interferometer are aligned parallel to one another, with a central chord passing through the center of the chamber, and the others in pairs, offset from the central by 12.75 cm and 25.5 cm respectively. Thus the chords measure the coarse radial density profile as well as provide a check against the spherical symmetry of the liner (if a perfectly spherical plasma liner is formed, the signals on the chord pairs should perfectly match). After passing through the plasma, the probe beams are compared with a reference beam of the interferometer in the Mach-Zender configuration, and the line-integrated electron density is computed based on the relative phase shift [21, 20].
@@ III.5 Spatially-Resolved Emission Spectroscopy
Spatially-resolved visible plasma emission spectra provides insight into LOS-integrated behavior of electron temperature via relative line emission intensities, and relative plasma density via observed Bremsstrahlung emission intensity. In the Multi-Chord Spectrometer (MCS), data is collected along 11 fiber-coupled lines of sight through 2.5 cm diameter collimators mounted outside the chamber. This light is sent through an imaging spectrometer, a Chromex 500is with 0.5 m focal length, a 1200 grooves/mm grating and 50 m slit for an effective resolution of 0.08 nm over the wavelength range used, 453.3 nm to 482.2 nm. A relative intensity calibration was applied based on a calibrated tungsten Thorlabs SLS201L lamp shining through the same set of windows and optics as the PLX plasma light traverses. An additional calibration measure was taken to normalize the total amount of light collected through each fiber; to do so, each fiber was illuminated individually with the lamp and imaged with constant camera settings, allowing for comparison of counts vs. wavelength for all fibers.

### 20 Feb 2024 v2

@@ Abstract
Plasma-jet-driven magneto-inertial fusion (PJMIF) is an alternative approach to controlled nuclear fusion which aims to utilize a line-replaceable dense plasma liner as a repetitive spherical compression driver. In this experiment, first measurements of the formation of a spherical Argon plasma liner formed from 36 discrete pulsed plasma jets are obtained on the Plasma Liner Experiment (PLX). Properties including liner uniformity and morphology, plasma density, temperature, and ram pressure are assessed as a function of time throughout the implosion process and indicate an apparent transition from initial kinetic inter-jet interpenetration to collisional regime near stagnation times, in accordance with theoretical expectation. A lack of primary shock structures between adjacent jets during flight implies that arbitrarily smooth liners may be formed by way of corresponding improvements in jet parameters and control. The measurements facilitate the benchmarking of computational models and understanding the scaling of plasma liners towards fusion-relevant energy density.
I Introduction
@@
Magneto-inertial fusion (MIF, also called magnetized target fusion or MTF) is an approach to controlled nuclear fusion which joins aspects of both magnetic confinement fusion and inertial confinement fusion methods, enabling controlled fusion at more modest values of both confinement time and density [1, 2, 3, 4, 5] than those in the purely inertial or magnetic confinement counterparts. Many MIF compression approaches and configurations have been proposed and studied, including the Magnetized Liner Inertial Fusion (MagLIF) Z-pinch configuration [6], shear-flow-stabilized Z pinches [7], spherical piston-driven liquid metal liners [8], rotationally stabilized liquid metal liners[9], and staged-gas-puff Z-pinches[10]. Notably, MagLIF experiments have successfully produced thermonuclear conditions and significant neutron yields from the compression of preheated magnetized deuterium fuel, on implosion timescales significantly longer than can be tolerated in pure inertial fusion.
@@
The plasma-jet-driven magneto-inertial fusion (PJMIF) concept is an MIF approach in which the requisite heavy compressive liner is formed from the merging and spherical convergence of an array of discrete plasma jets. PJMIF aims to access an attractive subsection of the MIF physics parameters space, namely that of the spherical, moderate velocity (50-150 km/s) compression of high-beta high-Hall-parameter targets which have shown promise in MagLIF [11, 12, 5]. If successful, this architecture would enable repetitive line-replaceable operation with a significant standoff distance between the fusion plasma and driver hardware. Achieving economical repetitive operation is a significant challenge facing all proposed pulsed fusion concepts, in particular those seeking to utilize electrical pulsed-power as the primary energy source [4, 11].
@@
There remain several key questions regarding the viability of the PJMIF approach, including whether conditions of sufficient energy density and fusion gain can be attained while maintaining the physical standoff that constitutes a key appeal of the approach, whether sufficient uniformity of a spherical liner can be attained from the merger of discrete jets in the face of jet-on-jet shock waves and Rayleigh-Taylor instabilities [13, 14, 15], and whether effective magnetic insulation of a standoff target plasma can be achieved in spherical implosion geometry [16]. Prior investigations have shed light on several of these questions, with analytical and computational modeling predicting the feasibility of significant fusion gain [17], the realization of high-performance modular plasma guns to produce controlled plasma jets [18], and laboratory experiments studying the merging and formation processes of multiple plasma jets at representative velocities [19, 20, 21, 22, 23]. Much of the experimental work has been done at the Plasma Liner Experiment (PLX) at Los Alamos, a moderate-scale experimental effort towards understanding the challenges and scaling of PJMIF. Early experimental campaigns at PLX focused on plasma gun development and generating suitable pulsed-power-driven jets [18, 24], while more recent campaigns have been aimed at understanding and characterizing jet merging behavior in obliquely merging jet pairs [19, 20], and liner subsection formation of up to 6 or 7 jets [4, 21].
@@
In the present work, fully spherical plasma liners have been formed and investigated at the Plasma Liner Experiment (PLX). This experiment uses 36 individual argon plasma jets, enabling observation of the integrated liner assembly process for the first time. The primary aim is to assess the uniformity and plasma characteristics of the full spherical liner formed by discrete jets and determine where future work must focus to close the gap between present liner quality and fusion-relevant characteristics. The discussion of this full liner experimental campaign is organized as follows: Section II details the pulsed power setup and plasma jet formation and tuning, Section III describes what plasma properties are measured and the diagnostics used to do so, and Section IV presents the results of the diagnostic suite and implications for PJMIF as a whole. Considerations necessary for continuation and an eventual reactor-scale device are explored in Section VI as well as a concluding summary of points.
II Experiment
@@
The PLX device consists of a 3-meter diameter chamber and 36 attached plasma guns, shown in Fig. 1 along with a rendered cross section of the chamber and notional plasma liner and target within [12]. The plasma guns are positioned quasi-spherically over the chamber, with an average full-angle between adjacent guns of 36 degrees and a minimum full-angle of 22 degrees due to the relative positioning of available gun attachment ports. Chamber vacuum base pressure is typically around 1 torr before plasma jets are injected into the chamber, increasing to 10 torr when a turbopump gate valve is closed before the shot.
@@
The plasma guns that supply the liner-forming plasma jets are contoured-gap coaxial plasma railguns - the contoured shape of the inner electrode alleviates the blowby instability in such geometries that can hinder outgoing jet quality. The guns used in the present PLX work are similar to those described in detail in Ref[18], although the current PLX ones are a more compact design. The stored energy per gun per pulse is 4.5-6 kJ, which accelerates argon plasma jets of total mass 1 mg to typical tunable velocities between 30 and 70 km/s out of a barrel that is 8.5 cm in diameter. Most of the plasma gun capacitor banks and spark-gap switches are mounted directly onto the chamber to minimize transfer inductance to the plasma load and increase energy coupling. Due to space access limitations around the lowest ports of the PLX chamber, nine guns are instead linked to their capacitor banks via an array of transmission lines, referred to hereafter as transmission line (TL) guns. Plasma properties of the jets at the time of ejection have been explored in previous publications [20]; typically, for a given jet around ejection time, eV and cm-3.
@@
Four distinct collections of capacitor banks are activated when firing a PLX shot. The first is a high voltage (19.0 kV) modified-Blumlein trigger network for the plasma gun gas valves. This sends 36 trigger pulses down coaxial cables to 36 spark gap switches that operate the gas valve capacitor banks of the plasma guns. Each plasma gun has a separate 100 J capacitor bank that actuates the gun’s fast gas puff valve, injecting neutral argon gas injection into the guns. An adjustable time delay passes while gas valves operate and neutral gas flows into the bore of the plasma guns, typically 480 microseconds in the current experiments. Thirdly, another modified-Blumlein trigger network fires, to breakdown the parallel spark-gaps for the main gun banks. These banks have 575 F rated to 5 kV peak voltage, and provide most of the current to form and accelerate the plasma jet. The main bank is discharged across a set of six spark-gap switches to the coaxial rail electrodes, again to minimize transfer inductance and prolong switch electrode lifetime. Operating voltages are set to 4.0 kV for non-TL guns, while the TL guns described above are charged to 4.5 kV to compensate for the increased transfer inductance of their transmission lines. Time is defined as when the main trigger pulse for the main accelerating banks is sent. Pressure adjustments can be made to many of the spark-gap switches on a per-gun basis to ensure they all break down properly with the trigger pulse. A resistive circuit in parallel with the pulse-forming-network (PFN) spark-gap trigger switches produces pre-ionization at the breech - this preionization reduces time jitter of the bulk plasma breakdown and subsequent acceleration generated by the main bank.
@@
Tuning the individual jet velocity is essential in this experiment, in order to ensure the jets arrive in the chamber at the same time and merge together with as high degree of spherical symmetry as possible. The component voltages described above can be adjusted for varying effects: a higher main trigger voltage can reduce firing jitter but increases probability of prefiring and arcing between components; main bank charge voltage establishes the overall energy with which the jets are ionized and accelerated; GV charge voltages establish how much gas is injected into the breach and subsequently accelerated. The most robust means of tuning jet velocity is adjusting the GV bank voltage and therefore the injected mass into which the accelerating PFN energy is deposited. To achieve a nominal jet front velocity of 50 km/s for each of the 36 jets, adjustments to this GV bank voltage for each individual gun are made approximately 2 seconds to firing to fine-tune the respective jet velocity. This tuning method enables significantly more flexible and easier adjustment than the previous method of introducing ballast impedence wires in series with the GV described in Appendix B of Ref[21]. Results in this paper are obtained using a nominal best tune of each gun to 50 km/s. The process of determining the best tune is described in Section III.2. Mach number , with , where adiabatic index is assumed and is the Argon ion mass. To estimate , we will use calculations from previous work examining the exiting jet for temperature and . With these values, estimated ejection Mach number in the present experiment is around .
@@
Main bank and GV discharges are monitored using a set of Rogowski coils on each gun: six on the main bank transmission lines and one on the GV transmission line. Signals from these are sent to a static integrator box to convert the measured to where is current. It has been empirically determined that jet ejection time corresponds to about 2.5 s after the peak of the main bank Rogowski current signal. Figure 2 shows a set of all 36 main bank Rogowski signals, the mean signal and standard deviation zone around it, and the mean ejection time given by the dashed line. Current ringing is observed after the primary accelerating current corresponding to the signal peak, as the underdamped capacitor circuit discharges its remaining energy into residual trailing plasma in the gun.
III Diagnostics
@@
A suite of diagnostics is used to interrogate the plasma liner properties throughout the merge and stagnation process. Time-resolved laser interferometry measurements of line-integrated plasma density are acquired over the duration of each shot, and a multi-frame fast camera images the plasma at several times throughout each shot. A collection of single-frame wide-view cameras are also used for imaging. Photodiode time-of-flight pairs are used measure individual jet velocities for each of the 36 plasma jets. Individual gun performance is also monitored using Rogowski coils that measure the main accelerating current of each gun (see Fig. 2). An imaging spectrometer provides electron temperature and charge state information and high-resolution spectrometer is used to verify electron density and jet slowing at chamber center. Diagnostics lacking inherent time-resolution are instead triggered at varying times for different shots to build up an average picture of temporal evolution of the liner. Figure 3 indicates generally the layout and viewing arrangements of most of the diagnostics in the experimental campaign (the wide-view cameras are not indicated, as each has a view of the entire chamber interior).
@@ III.1 Imaging
III.2 Photodiode Pairs
Each gun is equipped with two light-collecting fibers positioned along the jet trajectory and spaced by 2 cm which feed to two photodiodes having sufficiently fast risetime to track the jet (5 ns). The collection fibers are recessed into their holders by 6 mm to limit the light collection cone. The 2 cm separation distance was selected as a compromise between two competing considerations: there must be enough time separation between the signals that the available digitization rate of 40 MHz does not significantly contribute to uncertainty (i.e. there are sufficient samples separating the two signals), and the closer together the fibers are the signals are more likely to be closely correlated in shape which enables easier analysis.
Figure 4 shows a diagram of the two photodiode views of a jet, at two times of jet emergence from the gun cavity. The corresponding locations on normalized sample signal readouts are also indicated, along with the region within which a cross correlation is applied. The bounds of the cross correlation region are selected to be a few microseconds before the signal rises, and a few microseconds after the signal peak. This provides the jet front velocity. The jets consist of a distribution of directed velocities with the trailing portion being slower than the jet front, resulting in an extended column of plasma. The velocity analysis focuses on the front region because this the portion which forms the spherical liner to compress the MIF target and is therefore most relevant. The photodiode analysis process is automated and repeated for each gun and enables velocities to be measured immediately following each shot and tuned as desired by adjusting the GV bank voltage - this we call the tuned velocity setting.
This photodiode pair time-of-flight velocimetry methodology was cross-checked by taking a set of shots with a single gun at a time while positioning a single-tip Langmuir probe at chamber center, and obtaining a jet velocity via the plasma time-of-flight to the probe () and comparing to the velocity inferred via the photodiode signals (). A typical comparison is shown in Fig. 5. In this, we define the jet arrival time to the probe as the location corresponding to 20% of the full integral of the probe signal. The previously discussed jet ejection time is subtracted from this arrival time, and in conjunction with the travel distance being the PLX radius of 136.5 cm, provides . This generally matches well with the first peak-like feature in the signal, as seen in subfigure b.
@@ III.2 Photodiode Pairs
III.3 High Resolution Spectroscopy
@@ III.4 Multi-Chord Interferometry
Line-integrated electron density is obtained in this work via a multi-chord heterodyne laser interferometry system whose arrangement is laid out in Ref[26]. The laser emission is produced using a 320 mW, 561 nm solid state laser, is split into multiple lower-power beams and injected into the chamber through five single-mode fiber optic cables with each probe beam about 0.3 cm in diameter. The scene chords of the interferometer are aligned parallel to one another, with a central chord passing through the center of the chamber, and the others in pairs, offset from the central by 12.75 cm and 25.5 cm respectively. Thus the chords measure the coarse radial density profile as well as provide a check against the spherical symmetry of the liner (if a perfectly spherical plasma liner is formed, the signals on the chord pairs should perfectly match). After passing through the plasma, the probe beams are compared with a reference beam of the interferometer in the Mach-Zender configuration, and the line-integrated electron density is computed based on the relative phase shift [27, 26].
@@ III.5 Spatially-Resolved Emission Spectroscopy
Spatially-resolved visible plasma emission spectra provides insight into LOS-integrated behavior of electron temperature via relative line emission intensities, and relative plasma density via observed Bremsstrahlung emission intensity. In the Multi-Chord Spectrometer (MCS), data is collected along 11 fiber-coupled lines of sight through 2.5 cm diameter collimating lens of 6 cm focal length mounted outside the chamber, with waist size of 1 cm. This light is focused on the entrance slit of a Chromex 500is imaging spectrometer with 0.5 m focal length, a 1200 grooves/mm grating and 50 m slit for an effective resolution of 0.08 nm over the wavelength range used, 453.3 nm to 482.2 nm. A relative intensity calibration was applied based on a calibrated tungsten Thorlabs SLS201L lamp shining through the same set of windows and optics as the PLX plasma light traverses. An additional calibration measure was taken to normalize the total amount of light collected through each fiber; to do so, each fiber was illuminated individually with the lamp and imaged with constant camera settings, allowing for comparison of counts vs. wavelength for all fibers.